OpenStudy (anonymous):

For questions 1-2 multiply. State any excluded values? will give medal! 1. 2/2a x 2/a^2 2. q-3/-5 x 3q/q-4

OpenStudy (anonymous):

\[\frac{ 2 }{ 3a }\times \frac{ 2 }{ a ^{2} }\]